• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M Jones

    Log entry time 20:21:39 on February 22,1999

    Entry number 15293

    keyword=MCC fixed problem with cavity lock

    MCC fixed problem with cavity lock. From 16:30 until 20:00 we have had
    spotty beam with beam trips every 5-10 minutes. MCC found that the
    problem was a cavity on the South Linac that lost its lock and shift the
    beam energy by 10-3 and causing a ion chamber trips in Hall A.
    Hopefully should have smooth beam the rest of the night